# Initialization Schemes for Kolmogorov–Arnold Networks: An Empirical Study

### v1

@@ 4.3 Feynman Dataset Benchmarks
As a final benchmark, we turn to a subset of the Feynman dataset, restricted to dimensionless equations. The explicit formulas of the target functions and implementation details are provided in Appendix C.3. We evaluate the baseline, Glorot, and power-law initializations using the same “small” and “large” architectures defined in Section 4.2. Tables 3 and 4 report the results in terms of final training loss and relative $L^{2}$ error with respect to the reference solutions, for the small and large settings, respectively.
Table 3: Results on the Feynman benchmark for the “small” architecture ($G=5$, two hidden layers with 8 neurons each). Reported values correspond to median final training loss and relative $L^{2}$ error with respect to the reference solution. Best results per equation are shown in bold.
@@
Table 4: Results on the Feynman benchmark for the “large” architecture ($G=20$, three hidden layers with 32 neurons each). Reported values correspond to median final training loss and relative $L^{2}$ error with respect to the reference solution. Best results per equation are shown in bold.
@@ 5 Conclusion
In this work, we proposed and systematically evaluated new initialization strategies for spline-based KANs. Specifically, we introduced two theory-driven schemes inspired by LeCun and Glorot, including a variant with batch-normalized basis functions, as well as an empirical family of power-law initializations with tunable exponents. Through large-scale grid searches, training dynamics analysis, and evaluations on both PDE and function fitting benchmarks, we showed that initialization plays a crucial role in KAN performance. Our results demonstrate that while LeCun-inspired schemes offer limited benefits, Glorot-inspired initialization emerges as a strong candidate for parameter-rich architectures, and the empirical power-law family provides the most robust improvements overall, achieving faster convergence and lower errors across benchmarks. These findings highlight initialization as a key component of training KANs and identify effective practical strategies for the process.
@@ 5.1 Limitations and Future Work
The full code (including selected seeds for each experiment) and the processed data from the grid-search experiments are publicly available at https://github.com/srigas/KAN_Initialization_Schemes.
@@ Appendix D Indicative Results for Power-Law Grid-Search
Appendix E Neural Tangent Kernel Analysis
@@
E.1 NTK for PIML with RBA Weights
@@
E.2 Additional NTK Spectra
@@
For completeness, we report additional NTK spectra not included in the main text. Figures 1417 show the results for the remaining function fitting benchmarks ($f_{1}$, $f_{2}$, $f_{4}$, $f_{5}$), while Figures 18, 19 correspond to the Burgers’ and Helmholtz PDEs. All results are obtained using the “large” architecture ($G=20$, three hidden layers with 32 neurons each), consistent with the setting analyzed in Section 4.2.
@@
Figure 14: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{1}(x,y)$ under different initialization strategies.
@@
Figure 15: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{2}(x,y)$ under different initialization strategies.
@@
Figure 16: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{4}(x,y)$ under different initialization strategies.
@@
Figure 17: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{5}(x,y)$ under different initialization strategies.
@@
Figure 18: NTK eigenvalue spectra for the Burgers’ PDE benchmark under baseline, Glorot, and power-law initializations. Top row: spectra corresponding to the PDE residual term. Bottom row: spectra for the boundary/initial condition terms. Solid lines show the initialization, dashed lines show intermediate iterations, and dotted lines show the final iteration.
@@
Figure 19: NTK eigenvalue spectra for the Helmholtz PDE benchmark under baseline, Glorot, and power-law initializations. Top row: spectra corresponding to the PDE residual term. Bottom row: spectra for the boundary/initial condition terms. Solid lines show the initialization, dashed lines show intermediate iterations, and dotted lines show the final iteration.

### v2

@@ 4.3 Feynman Dataset Benchmarks
As a final benchmark, we turn to a subset of the Feynman dataset, restricted to dimensionless equations. The explicit formulas of the target functions and implementation details are provided in Appendix C.3. Having already established via empirical evidence and NTK analysis that the LeCun-based schemes are consistently underperforming, we benchmark here only the two competitive initializations (Glorot and power-law) alongside the baseline. We again fix the power-law exponents to $(\alpha,\beta)=(0.25,1.75)$ and evaluate all methods using the same “small” and “large” architectures defined in Section 4.2. Tables 3 and 4 report the results in terms of final training loss and relative $L^{2}$ error with respect to the reference solutions, for the small and large settings, respectively.
Table 3: Results on the Feynman benchmark for the “small” architecture ($G=5$, two hidden layers with 8 neurons each). Reported values correspond to median final training loss and relative $L^{2}$ error with respect to the reference solution. Best results per equation are shown in bold. The power-law initialization uses $\alpha=0.25,\beta=1.75$.
@@
Table 4: Results on the Feynman benchmark for the “large” architecture ($G=20$, three hidden layers with 32 neurons each). Reported values correspond to median final training loss and relative $L^{2}$ error with respect to the reference solution. Best results per equation are shown in bold. The power-law initialization uses $\alpha=0.25,\beta=1.75$.
@@ 5 Conclusion
In this work, we proposed and systematically evaluated new initialization strategies for spline-based KANs. Specifically, we introduced two theory-driven schemes inspired by LeCun and Glorot, including a variant with normalized basis functions, as well as an empirical family of power-law initializations. Through large-scale grid searches, we identified favorable exponent ranges for the power-law method for both function fitting and forward PDE problems. Across all evaluations, including loss curve analysis and NTK dynamics, we showed that initialization plays a critical role in KAN performance. In particular, our results demonstrate that while LeCun-inspired schemes offer limited benefits, Glorot-inspired initialization emerges as a strong candidate for parameter-rich architectures, and the empirical power-law family provides the most robust improvements overall, achieving faster convergence and lower errors across benchmarks. These findings highlight initialization as a key component of training KANs and identify effective practical strategies for the process.
@@ 5.1 Limitations and Future Work
The full code (including selected seeds for each experiment) and the processed data from the grid-search experiments are publicly available at the following GitHub repository: https://github.com/srigas/KAN_Initialization_Schemes.
Acknowledgments
S. R. and G. A. are supported by the European Union’s Horizon Europe Framework Programme for Research and Innovation, under the PATH project (Grant Agreement No. 101253520). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union, which cannot be held responsible for them.
@@ Appendix D Indicative Results for Power-Law Grid-Search
Appendix E Training Curves with Learning Rate Scheduling
In Section 4.2 of the main text, the training curves shown in Figures 1 and 2 were obtained using a fixed learning rate in order to isolate the effect of initialization, as initialization and learning-rate adaptability are known to interact (e.g., [Yang et al., 2021]). However, the fixed learning rate induces oscillations in the loss curves, particularly for the Glorot and power-law schemes in the larger architectures. To verify that these oscillations are purely an artifact of the constant learning rate, we repeat the same training experiments using a learning-rate scheduler.
Figure 14: Training loss curves for function fitting benchmarks under baseline, LeCun-numerical, Glorot and power-law ($\alpha=0.25,\beta=1.75$) initializations when using a learning-rate scheduler. Results are averaged over five seeds, with shaded regions indicating the standard error. Top row: “small” architecture ($G=5$, two hidden layers with 8 neurons each). Bottom row: “large” architecture ($G=20$, three hidden layers with 32 neurons each).
Figure 15: Training loss curves for forward PDE benchmarks under baseline, LeCun-numerical, Glorot and power-law ($\alpha=0.25,\beta=1.75$) initializations when using a learning-rate scheduler. Results are averaged over five seeds, with shaded regions indicating the standard error. Top row: “small” architecture ($G=5$, two hidden layers with 8 neurons each). Bottom row: “large” architecture ($G=20$, three hidden layers with 32 neurons each).
Figure 14 shows the results for the function fitting benchmarks, where all settings are identical to those in the main text except for the use of a learning-rate scheduler: training begins with a learning rate of $10^{-3}$, followed by exponential decay with decay factor $0.9$ every 50 iterations. Similarly, Figure 15 shows the results for the PDE benchmarks, again using all the same hyperparameters as in the main text except for the scheduler: training begins with a learning rate of $10^{-3}$, followed by exponential decay with decay factor $0.85$ every 100 iterations.
Across both sets of benchmarks, the learning-rate schedulers eliminate the oscillatory behavior observed under a fixed learning rate, yielding smoother training curves. Quantitatively, the final losses are slightly higher than those reported in the main text, due to the learning rate decaying even in regions where a larger step size would allow for further progress. Nonetheless, this influences only the numerical values: the qualitative picture remains unchanged, and the relative performance ordering of the initialization schemes is consistent with the fixed learning rate setting.
Appendix F Neural Tangent Kernel Analysis
@@
F.1 NTK for PIML with RBA Weights
@@
F.2 NTK Spectra for Varying Power-Law Exponents
To complement the heatmaps of Appendix D and to further illustrate the robustness of the power-law initialization, we examine in this Appendix how the NTK spectrum varies across different $(\alpha,\beta)$ configurations. The goal of this analysis is twofold. First, it provides an NTK-based view of the “good regions” identified in the grid search, showing how favorable exponent choices correspond to well-conditioned and stable kernels. Second, it demonstrates that the power-law scheme is not sensitive to a single finely tuned pair of exponents: more than one $(\alpha,\beta)$ combinations within the identified range yield well-behaved spectra. This supports the idea that one may tune the exponents once per problem domain and thereafter select any configuration from the favorable region.
Figure 16 displays the NTK eigenvalue spectra for all exponent pairs considered in the grid search, using the “large” architecture ($G=20$, three hidden layers with 32 neurons each) and the function fitting target $f_{3}(x,y)$. Figure 17 shows the corresponding results for the PDE residual term of Burgers’ equation. In both cases, well-conditioned spectra concentrate in the same regions suggested by the grid-search results.
Figure 16: NTK eigenvalue spectra for the large architecture ($G=20$, three hidden layers with 32 neurons each) on the function fitting target $f_{3}(x,y)$, shown for all $(\alpha,\beta)$ configurations considered in the grid search. Each panel corresponds to one exponent pair and displays spectra at initialization, mid-training, and convergence.
a
Figure 17: NTK eigenvalue spectra for the large architecture ($G=20$, three hidden layers with 32 neurons each) on the PDE residual term of Burgers’ equation, shown for all $(\alpha,\beta)$ configurations considered in the grid search. Each panel corresponds to one exponent pair and displays spectra at initialization, mid-training, and convergence.
F.3 Additional NTK Spectra
@@
For completeness, we report additional NTK spectra not included in the main text. Figures 1821 show the results for the remaining function fitting benchmarks ($f_{1}$, $f_{2}$, $f_{4}$, $f_{5}$), while Figures 22, 23 correspond to the Burgers’ and Helmholtz PDEs. All results are obtained using the “large” architecture ($G=20$, three hidden layers with 32 neurons each) and values $\alpha=0.25$, $\beta=1.75$ for the power-law initialization, consistent with the setting analyzed in Section 4.2.
@@
Figure 18: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{1}(x,y)$ under different initialization strategies. Results correspond to the “large” architecture ($G=20$, three hidden layers with 32 neurons each). The power-law initialization uses $\alpha=0.25,\beta=1.75$.
@@
Figure 19: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{2}(x,y)$ under different initialization strategies. Results correspond to the “large” architecture ($G=20$, three hidden layers with 32 neurons each). The power-law initialization uses $\alpha=0.25,\beta=1.75$.
@@
Figure 20: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{4}(x,y)$ under different initialization strategies. Results correspond to the “large” architecture ($G=20$, three hidden layers with 32 neurons each). The power-law initialization uses $\alpha=0.25,\beta=1.75$.
@@
Figure 21: Eigenvalue spectra of the NTK matrix at initialization (solid blue), intermediate iterations (dashed teal), and final iteration (dashed green) for function fitting benchmark $f_{5}(x,y)$ under different initialization strategies. Results correspond to the “large” architecture ($G=20$, three hidden layers with 32 neurons each). The power-law initialization uses $\alpha=0.25,\beta=1.75$.
@@
Figure 22: NTK eigenvalue spectra for the Burgers’ PDE benchmark under baseline, LeCun-numerical, Glorot, and power-law ($\alpha=0.25,\beta=1.75$) initializations. Top row: spectra corresponding to the PDE residual term. Bottom row: spectra for the boundary/initial condition terms. Solid blue lines show the initialization, dashed teal lines show intermediate iterations, and dashed green lines show the final iteration.
@@
Figure 23: NTK eigenvalue spectra for the Helmholtz PDE benchmark under baseline, LeCun-numerical, Glorot, and power-law ($\alpha=0.25,\beta=1.75$) initializations. Top row: spectra corresponding to the PDE residual term. Bottom row: spectra for the boundary/initial condition terms. Solid blue lines show the initialization, dashed teal lines show intermediate iterations, and dashed green lines show the final iteration.